OpenStudy (alibaby99):

A car is designed to last an average of 12 years with a standard deviation of 0.8 years. What is the probability that a car will last less than 10 years? 0.621% 6.21% 93.79% 99.379%

OpenStudy (will.h):

10-12/0.8=-2.5 P(-2.5)=0.0062 0.62%

OpenStudy (mathmale):

Borrowing from Will.H: "10-12/0.8=-2.5" is the z-score here, and is appropriate. How would you answer the question using this z-score? z=(10-12)/0.8=-2.5

OpenStudy (mathmale):

"Less than 10 years" here means anything from 0 to just under 10 years.